OpenStudy (anonymous):

does anyone know how to solve: ^-3 x ^-128. and show how you obtain answer.

OpenStudy (anonymous):

^-3 x ^-128 ?!!!? It's new to me! Coud you just write it on the board?

OpenStudy (anonymous):

\[ \sqrt{-3} \times \sqrt{-128}\]

OpenStudy (anonymous):

= i √3 * i √128 = - √3 * √ 2^7 = -8 √6

OpenStudy (anonymous):

ok.I got \[\sqrt{3} \pm \sqrt{2}\]

OpenStudy (anonymous):

how did you get your answer? please and thanks.

OpenStudy (anonymous):

Which part you don't understand?

OpenStudy (anonymous):

pretty much none of it. did you multiply the numbers 3and2 under the radicand to get 6?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

√ - 128 = √-1 * √ 2^7 = i * √2^6 * 2 = 2^3 i √2 = 8i √2

OpenStudy (anonymous):

Is it clear now?

OpenStudy (anonymous):

oh yes, it is. thanks.

OpenStudy (anonymous):

No more question?

OpenStudy (anonymous):

nope. I'm all set.

OpenStudy (anonymous):

That's what makes my day :)